OpenStudy (anonymous):

the dimensions of a rectangular solid are in the ratio of 1:6:6 and its volume is 562.5 cu. cm. find the volume of a cube whose surface area is equal to the surface area of a rectangular solid

OpenStudy (anonymous):

let the dimensions be x,6x,6x,then x*6x*6x=562.5 x^3=562.5/36=15.625 x=2.5 sides are 2.5,15,15 surface area s=2(l*b+b*h+l*h) =2(2.5*15+15*15+2.5*15) =2*15(2.5+15+2.5) =30*20=600 Let y be the edge of cube ,then surface area of cube S=6y^2=600 y^2=600/6=100 y=sqrt100=10 volume of cube=10^3=1000 cm^3